The common share of ABC, corp., is selling for $42 a share and investors require a 15% return on the stock. If two thirds of the return on this stock are derived from dividends yield and the other one third is derived from capital gain yield. What is the amount of the current dividend D0?

Question 27 options:

1)

3.65

2)

$3.86

3)

$4.00

4)

$4.15
Answer #1

Capital Gain =15%/3 =5%
Price in year before =42/(1+5%) =40

Dividend Yield =15%*2/3 =10%
Current Dividend =Dividend Yield*Price a year ago =10%*40 =4